OpenStudy (anonymous):

1.Are 24, 45, and 51 the side measurements of a special right triangle?

Directrix (directrix):

24 = 8 *3 45 = 15*3 51 = 17*3 ---------------- 8, 15, 17 is a Pythagorean Triple. Therefore, 24, 45, and 51 are sides of a right triangle.

OpenStudy (anonymous):

thank u

Directrix (directrix):

Glad to help.